OpenStudy (anonymous):

MEDAL REWARDED!! Penny's savings account pays an interest rate of 3%. If her balance is $10,100 at the end of the year , what will her balance be including the interest she earned? A. $10,403.00 B. $13,130.00 C. $23,300.00 D. $30,300.00

OpenStudy (anonymous):

wouldnt my answer be D?

OpenStudy (anonymous):

@e.cociuba

OpenStudy (jdoe0001):

hmm assuming the interest rate pays annually she gets her balance PLUS her interest

OpenStudy (anonymous):

yea so wouldnt that make my answer D?

OpenStudy (jdoe0001):

let's see D 30,300.00 - 10,100 = 20,200 20,200 is twice as much as 10,100, that'be an interest rate of twice as much, or a whooping 200% heck, if a bank pays that much, sign me in

OpenStudy (anonymous):

so is D correct or no lo?

OpenStudy (jdoe0001):

D has an interest rate of 200%, no 3% so..... no dice on that one her balance plus her interest what's her interest? well, 3% of 10,000, \(\bf 3\%\textit{ of } 10,000\implies \cfrac{3}{100}\times 10,100\)

OpenStudy (jdoe0001):

mm... have a typo there, anyhow \(\bf 3\%\textit{ of } 10,100\implies \cfrac{3}{100}\times 10,100\)

OpenStudy (anonymous):

i get 303

OpenStudy (jdoe0001):

yeap, that's her earned interest

OpenStudy (anonymous):

okay so now what? how do i find my answer than?

OpenStudy (jdoe0001):

well.. her balance plus THAT

OpenStudy (anonymous):

oh it would be A!

OpenStudy (jdoe0001):

:)

OpenStudy (jdoe0001):

10,100 + 303 = 10,403